May 18, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, May 18, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. GlueX Detector Performance Document
  4. Data Validation
  5. Calibration Train


  1. Announcements
  2. Calibrations for REST Production
  3. Subdetector Reports
    • Calorimetry
    • Tracking
    • PID
    • Photon Beam
  4. AOB


Attending: Sean (NU); Mark I., Beni, Adesh, Alex B., Alexander A., Nathan, Justin, Mark D., Elton (JLab); Curtis, Naomi, Mike S., Will M., Luke K. (CMU); Matt S. (IU); Maria (MIT)

  1. Announcements
    • The batch farm was paused for a day or so for the transition to a new writeable cache disk, which is where our REST data is kept and where raw data files are staged. Mark Ito discussed the situation and said that we should think of this system as a filesystem-based interface to the tape system. We have a total quota of 115 TB on the lustre disks, spread out among /cache, /volatile, and /work. Each disk has its own individual quotas as well.
    • The final calibration run is ~80% done, should be enough for most calibrations that need to be finished. These runs include our "high priority" runs.
  2. We reviewed the status of the calibrations for REST production planned at the end of the month.
    • Simon is working on some geometry and tracking issues. He is away at the Baryon conference, will check in with him on Monday.
    • Mike S. will update some CDC calibration parameters, and is on track.
    • Beni gave an overview of the background of the TOF TDC shift parameter, which resolves a 6-fold clock ambiguity and looks to be the cause of the TOF matching inefficiency reported at the Collaboration Meeting. He has been looking into the issue in depth and noted that it seems like this parameter, which changed at the beginning of the run to a value of 1, changed during the "bracketgate" fix downtime to a value of zero. This corresponds to a change in the TS firmware. He is currently testing a change in how this value is applied, where due to differing conventions, a '<' should be '<='. It is unclear how the parameter values are actually changing, and if this fix works, why the data from last spring looks fine.
    • Will has started the BCAL gain calibrations on the high priority runs, and is working to study the systematics on the pi0 mass. After finishing these calibrations, he plans to improve the clustering algorithm for charged particles.
    • Mark Dalton reported that work to include George's new effective velocities in the DBCALPoint objects will be done by production launch.
    • Elton asked about possibly including Ahmed's BCAL LED monitoring plugin. Sean suggested that some more trigger types in the data could be skimmed out and will discuss this with with Paul and Justin.
    • Alex B. is almost done with the TAGM timewalks, and is just investigating some inconsistencies that popped up. He will then nail down the software thresholds.
    • Adesh is busy with the FCAL gain calibrations for this last data, and is looking at the LED data to try to understand the drifts in the pi0 peak.
      • Matt asked about the possibility of applying some simple scale and non-linear corrections at analysis time on REST data, ala CLEO. The BCAL group is interested in this as well. Such a scheme could be accomodated, and the discussion of the details will take place offline.
    • Nathan ran into some behavior in the TAGH hit merging that is more complicated than expected. He is shifting his focus to updating the timing calibrations.
    • Beni pointed out that the z position of the TOF planes is swapped in the current geometry. This won't have a large effect on reconstruction, but the calibrations could change. The plan is to fix this by the end of the month as well.
  3. Simulations
    • Mark Ito asked if new simulations are needed soon. A "small" sample with the current reconstruction code would be useful for data/MC matching, since enough has changed since the original sim1. A small set could be generated at JLab before the data launch if there is enough headroom. It could also be generated offsite. Sean will start generating condition files, and the details will be discussed at next week's Offline meeting.